### v2

# AI-Enhanced Real-Time Wi-Fi Sensing Through Single Transceiver Pair

Abstract
The advancement of next-generation Wi-Fi technology heavily relies on sensing capabilities, which play a pivotal role in enabling sophisticated applications. In response to the growing demand for large-scale deployments, contemporary Wi-Fi sensing systems strive to achieve high-precision perception while maintaining minimal bandwidth consumption and antenna count requirements. Remarkably, various AI-driven perception technologies have demonstrated the ability to surpass the traditional resolution limitations imposed by radar theory. However, the theoretical underpinnings of this phenomenon have not been thoroughly investigated in existing research. In this study, we found that under hardware-constrained conditions, the performance gains brought by AI to Wi-Fi sensing systems primarily originate from two aspects: prior information and temporal correlation. Prior information enables the AI to generate plausible details based on vague input, while temporal correlation helps reduce the upper bound of sensing error. Building on these insights, we developed a real-time, AI-based Wi-Fi sensing and visualization system using a single transceiver pair, and designed experiments focusing on human pose estimation and indoor localization. The system operates in real time on commodity hardware, and experimental results confirm our theoretical findings.
publicationid: pubid:
I Introduction
Sensing plays a crucial role in the evolution of next-generation Wi-Fi technology [1]. With its high penetration rate and diverse non-contact sensing functionalities, device-free Wi-Fi sensing has presented extensive application potentials in areas such as home security, abnormal behavior detection, and elderly or child care [2].
Despite the significant potentials of device-free Wi-Fi sensing technologies, their large-scale implementation remains challenging. The received signal strength indicator (RSSI) [3, 4] or channel state information (CSI) [5] of a target at different positions in the environment can be used as fingerprints. Researchers construct a fingerprint database for indexing and compare received fingerprint data with the stored entries. The spatial coordinates corresponding to the closest match in the database are then used as the estimated positioning result. These methods can only be used in locations where large amounts of data have been collected in advance. Another approach employs radar technology for sensing, with inverse synthetic aperture radar (ISAR) [6] and multi-signal classification (MUSIC) [7] along with its derived algorithms [8] being typical representative methods in this category. These technologies often require a large number of antennas and wide bandwidth that are difficult to implement with commodity Wi-Fi devices.
Some studies have proposed device-free Wi-Fi sensing technology suitable for large-scale applications. For instance, Widar2.0 [9] requires at least a one antenna transmitter and a three antenna receiver to facilitate positioning; however, its limited system resolution hampers more precise sensing applications such as human pose estimation (HPE) or breathing detection.
Recently, methods leveraging artificial intelligence (AI) have garned increasingly attention, demonstrating superior resolution and accuracy compared to traditional technologies. Most of these methods leverage multiple transceiver units to enhance accuracy [10, 11], while others focus on achieving high-precision sensing with minimal hardware [12], demonstrating notable success. For instance, CSI2Depth [13] utilizes the Wi-Fi CSI data from the MM-Fi dataset [14], generating depth images based on CSI collected by a pair of TP-Link N750 Wi-Fi APs.
Notably, some AI-based methods have demonstrated precision that significantly exceeds the theoretical resolution limit in radar theory (see Section II). Although certain studies suggest that the Fresnel zone model [15] can explain the high accuracy of Wi-Fi sensing, when only a single transmitter-receiver pair is used, this model only accounts for the system’s high precision in the direction perpendicular to the line connecting the transceivers. According to this model, additional devices must be deployed to achieve similar accuracy in other directions. [16] reveals the influence of target motion on CSI amplitude, providing a theoretical basis for motion pattern recognition based on CSI. However, it still fails to explain why AI-based methods also achieve remarkable performance in regression tasks such as HPE. To the best of our knowledge, there is currently no widely accepted theoretical framework that can fully explain where the performance gains brought by AI to Wi-Fi sensing originate.
In this paper, we first propose a theoretical explanation for the performance gains achieved by AI. Then, based on our theoretical insight, we develop an AI-based real-time Wi-Fi sensing and visualization system and conduct experiments to validate the proposed theoretical explanation. The main contributions of this work are as follows:
  • We suggest that the performance improvement brought by AI to regression tasks fundamentally stems from prior information and temporal correlation. First, by leveraging prior knowledge about the target structure acquired during training, AI can generate detailed target descriptions that surpass the resolution limits imposed by the system’s radar aperture. Second, AI-based methods can effectively utilize temporal dependencies in the data, thereby narrowing the space of plausible estimation results and enhancing perceptual accuracy.
  • Based on our theoretical findings, we constructed an AI-based Wi-Fi sensing system using only a single transmitter-receiver pair. The system achieves an average localization error of $0.6124\text{ m}$ and an average HPE error of $0.2189\text{ m}$. Through experiments, we have confirmed the performance improvement brought by temporal correlation. Furthermore, we have observed indirect evidence of AI leveraging prior information.
  • We further optimized the data processing pipeline, thus enabling the Wi-Fi sensing system to operate in real time with simultaneous result visualization. The system achieves a frame rate of no less than $42\text{ fps}$ on commodity hardware.
The remainder of this paper is organized as follows. Section II presents the theoretical analysis of the performance gains achieved through AI. Section III introduces the design and implementation of the proposed Wi-Fi sensing system. Section IV details the experimental setup and results. Section presents the architecture of the real-time sensing and visualization system. Finally, Section VI concludes the paper.
II Theoretical Analysis of the Performance Gains
II-A Spatial Resolution Analysis Based on Radar Theory
From the perspective of traditional radar theory, an analysis of a typical Wi-Fi sensing system with only a single transceiver pair reveals that its spatial resolution is limited to the meter level under such a constrained hardware setup. We consider a general scenario in which the distance between the transmitter and receiver is arbitrary. The analysis of spatial resolution can be divided into two aspects: angular resolution and distance resolution.
Figure 1: Angle-of-arrival estimation in a typical Wi-Fi sensing system.
For angle-of-arrival (AoA) estimation, consider a signal transmitted from the transmitting antennas, reflected by a target, and finally received by the receiving antennas, as shown in Fig. 1, where the angle-of-departure is $\theta_{\text{t}}$, and the AoA is $\theta_{\text{r}}$. The distance between adjacent transmitting antennas is denoted as $d_{\text{t}}$, and the distance between adjacent receiving antennas is denoted as $d_{\text{r}}$. Denote the signal from transmitting antenna $m$ to the receiving antenna $n$ as $\text{signal}_{mn}$, then the phase difference between $\text{signal}_{mn}$ and $\text{signal}_{00}$ is
$\Delta\phi_{mn}=\frac{2\pi}{\lambda}(md_{\text{t}}\sin\theta_{\text{t}}+nd_{\text{r}}\sin\theta_{\text{r}}),$ (1)
where $\lambda$ is the wavelength of the transmitted signal. Let the transmitted signal be $x(t)$, then the received signal can be represented as
$y(t)=\sum_{m=0}^{m=N_{\text{t}}-1}\sum_{n=0}^{n=N_{\text{r}}-1}x(t)e^{j\Delta\phi_{mn}},$ (2)
which can be regarded as a 2-D Fourier transform, so the angular resolution of the established system can be regarded as the resolution of this transform in the direction of $\theta_{\text{r}}$. Thus, the angular estimation resolution of the system is
$\Delta\theta_{\text{r}}=\frac{\lambda}{N_{\text{r}}d_{\text{r}}\left\lvert\cos\theta_{\text{r}}\right\rvert}$. (3)
Under typical Wi-Fi sensing system configurations, $d_{r}=\frac{\lambda}{2}$ and $N_{\text{r}}=3$, then the resulting angular resolution is $\Delta\theta_{\text{r}}=0.667\ \text{rad}$.
For distance estimation, time-of-flight (ToF) estimation can only determine an ellipse with the transmitter and receiver as its foci, within which the target may be located. Distance estimation from the target to the receiver must therefore be combined with AoA estimation results. Since
$\text{d}l_{\text{t}}=\frac{l_{\text{r}}-d_{\text{rt}}\cos\theta_{\text{r}}}{l_{\text{t}}}\text{d}l_{\text{r}},$ (4)
where $l_{\text{t}}$ is the distance between the target and the transmitting antennas, $l_{\text{r}}$ is the distance between the target and the receiving antennas, $d_{\text{rt}}$ is the distance between the transmitting and receiving antennas. Let $l_{\text{t}}+l_{\text{r}}=L$, $L>d_{\text{rt}}$, then
$\text{d}l_{\text{t}}=\left(1-\frac{2d_{\text{rt}}^{2}\sin^{2}\theta_{\text{r}}}{d_{\text{rt}}^{2}+L^{2}-2d_{\text{rt}}L\cos\theta_{\text{r}}}\right)\text{d}l_{\text{r}}.$ (5)
According to radar theory, $L$ is estimated to have an upper resolution of
$\Delta L=\frac{c}{B},$ (6)
where $c$ is the speed of light, and $B$ is the signal bandwidth. Since $\Delta l_{\text{t}}=\Delta L-\Delta l_{\text{r}}$, substituting $\Delta l_{\text{t}}$ and $\Delta l_{\text{r}}$ for $\text{d}l_{\text{t}}$ and $\text{d}l_{\text{r}}$, then (5) can be transformed into
$\Delta l_{\text{r}}=\frac{d_{\text{rt}}^{2}+L^{2}-2d_{\text{rt}}L\cos\theta_{\text{r}}}{2(L-d_{\text{rt}}\cos\theta_{\text{r}})^{2}}\frac{c}{B}.$ (7)
When $\theta_{\text{r}}=0$, $\Delta l_{\text{r}}$ reaches its minimum value. For typical Wi-Fi sensing systems, the bandwidth ranges from $10\sim 40\ \text{MHz}$ [17]. Therefore, from (7), the upper distance estimation resolution is $\Delta l_{\text{r}}=3.75\ \text{m}$. Then the total spatial resolution is
$\Delta l=\sqrt{\Delta l_{\tau}^{2}+\left(2l_{\text{r}}\tan\frac{\Delta\theta_{\text{r}}}{2}\right)^{2}}.$ (8)
Since $\Delta l\geq\Delta l_{\text{r}}$, the resolution is not enough for accurate sensing tasks such as HPE.
II-B Performance Gain Brought by Prior Information
Figure 2: An example of the collected CSI. The area within the red box exhibits clear sinusoidal characteristics.
We use indoor HPE as an example for analysis. Treat CSI as the sum of paths affected only by static objects and paths influenced by dynamic objects, we have
$h(f,t)=e^{-j\phi(f,t)}\left(h_{\text{s}}(f)+\sum_{l=1}^{L_{\text{d}}}\alpha_{l}(f,t)e^{-j2\pi\frac{d_{l}(t)f}{c}}\right),$ (9)
where $f$ denotes the carrier frequency, $t$ represents time, $\phi(f,t)$ is the phase distortion caused by thermal noise and clock asynchronism at the receiver and transmitter, $h_{\text{s}}(f)$ denotes the CSI of paths affected only by static objects, $L_{\text{d}}$ is the number of paths influenced by moving objects, $\alpha_{l}(f,t)$ represents the attenuation of the $l$-th path, $d_{l}(t)$ denotes the length of the $l$-th path, and $c$ is the speed of light. According CSI-Speed model,
$\|h(f,t)\|^{2}=$ $\ \|h_{\text{s}}(f)\|^{2}$ (10)
$+2\sum_{l=1}^{L_{\text{d}}}\|h_{\text{s}}(f)\alpha_{l}(f,t)\|$
$\quad\times\cos\left[2\pi\left(\int_{0}^{t}\frac{v_{l}(\tau)}{c}f\,\mathrm{d}\tau+\frac{d_{l}(0)f}{c}\right)+\varphi_{sl}\right]$
$+\sum_{k=1}^{L_{\text{d}}}\sum_{l=1}^{L_{\text{d}}}\|\alpha_{k}(f,t)\alpha_{l}(f,t)\|$
$\quad\times\cos\left[2\pi\frac{(d_{k}(t)-d_{l}(t))f}{c}+\varphi_{kl}\right],$
where $v_{l}(\tau)$ denotes the rate of change in the length of path $l$, and $\varphi_{sl}$, $\varphi_{kl}$ represent constant phase shifts introduced by reflection. In indoor environments, due to strong reflections from walls, $\|h_{\text{s}}(f)\|\gg\|\alpha_{l}(f,t)\|$, thus the third term in (10) can be neglected. Therefore, the CSI in indoor environments can be modeled as a superposition of a static component and multiple dynamic sinusoidal components. This sinusoidal behavior is clearly observed in the actually collected CSI data over time, as demonstrated in Fig. 2.
Consequently, different velocity combinations manifest in the CSI as distinct superpositions of sinusoidal signals. The characteristics of these superimposed sinusoidal signals can be utilized to recognize the target’s actions, forming the theoretical basis for AI-based action classification tasks.
In regression tasks, these features can still yield performance gains. For HPE, the AI does not need to determine the precise coordinates of each joint. Instead, it can generate a human pose based on prior knowledge of human anatomy learned during training. We refer to this gain as the performance gain brought by prior information. A similar form of gain is also observed in related fields, such as deep learning-based super-resolution in image processing. Subsequent experiments will present an indirect evidence supporting this claim (as discussed in Section IV. D.).
II-C Performance Gain Brought by Temporal Correlation
Figure 3: CSI based Wi-Fi sensing process.
[…]
$\begin{split}J_{(d,\varepsilon)}^{(n)}=&\{(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})\in\mathcal{M}^{n}\times{\widehat{\mathcal{M}}}^{n}:\\ &\left\lvert-\frac{1}{n}\log p(\bm{m}_{i}^{n})-H(\bm{M}_{i})\right\rvert<\varepsilon,\\ &\left\lvert-\frac{1}{n}\log p(\widehat{\bm{m}_{i}}^{n})-H(\widehat{\bm{M}_{i}})\right\rvert<\varepsilon,\\ &\left\lvert-\frac{1}{n}\log p(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})-H(\bm{M}_{i},\widehat{\bm{M}_{i}})\right\rvert<\varepsilon,\\ &\left\lvert d(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})-E[d(\bm{M}_{i},\widehat{\bm{M}_{i}})]\right\rvert<\varepsilon\},\end{split}$ (11)
where $H(\cdot)$ represents information entropy, $\varepsilon>0$, $d(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})=\frac{1}{n}\sum_{j=1}^{n}d(\bm{m}^{(j)}_{i},\widehat{\bm{m}_{i}}^{(j)})$, and $E[\cdot]$ represents mathematical expectation. The pair $(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})\in J_{(d,\varepsilon)}^{(n)}$ can also be described as $\bm{m}_{i}^{n}$ and $\widehat{\bm{m}_{i}}^{n}$ are distortion typical.
Lemma 1. If $(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})\in J_{(d,\varepsilon)}^{(n)}$, then
$p(\widehat{\bm{m}_{i}}^{n}|\bm{m}_{i}^{n})\leq p(\widehat{\bm{m}_{i}}^{n})2^{n[I(\widehat{\bm{M}_{i}};\bm{M}_{i})+3\varepsilon]},$ (12)
where $I(\cdot)$ represents mutual information, $\varepsilon>0$.
Proof:
$\begin{split}p(\widehat{\bm{m}_{i}}^{n}|\bm{m}_{i}^{n})&=\frac{p(\widehat{\bm{m}_{i}}^{n},\bm{m}_{i}^{n})}{p(\bm{m}_{i}^{n})}=p(\widehat{\bm{m}_{i}}^{n})\frac{p(\widehat{\bm{m}_{i}}^{n},\bm{m}_{i}^{n})}{p(\widehat{\bm{m}_{i}}^{n})p(\bm{m}_{i}^{n})},\\ &\leq p(\widehat{\bm{m}_{i}}^{n})\frac{2^{-n[H(\widehat{\bm{M}_{i}},\bm{M}_{i})-\varepsilon]}}{2^{-n[H(\widehat{\bm{M}_{i}})+\varepsilon]}2^{-n[H(\bm{M}_{i})+\varepsilon]}},\\ &=p(\widehat{\bm{m}_{i}}^{n})2^{n[I(\widehat{\bm{M}_{i}};\bm{M}_{i})+3\varepsilon]}.\end{split}$ (13)
[…]
$\begin{split}J_{\text{t}(d,\varepsilon)}^{(n)}=&\{(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})\in\mathcal{M}^{n}\times{\widehat{\mathcal{M}}}^{n}:\\ &\left\lvert-\frac{1}{n}\log p(\bm{m}_{i}^{n})-H(\bm{M}_{i})\right\rvert<\varepsilon,\\ &\left\lvert-\frac{1}{n}\log p(\widehat{\bm{m}_{i}}^{n}|\widehat{\bm{h}}_{i-1,\cdots,0})-H(\widehat{\bm{M}_{i}}|\widehat{\bm{H}}_{i-1,\cdots,0})\right\rvert<\varepsilon,\\ &\left\lvert-\frac{1}{n}\log p(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})-H(\bm{M}_{i},\widehat{\bm{M}_{i}})\right\rvert<\varepsilon,\\ &\left\lvert d(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})-E[d(\bm{M}_{i},\widehat{\bm{M}_{i}})]\right\rvert<\varepsilon\},\end{split}$ (14)
where $\varepsilon>0$, $\widehat{\bm{h}}_{i-1,\cdots,0}$ means $\widehat{\bm{h}_{i-1}},\widehat{\bm{h}_{i-2}},\cdots,\widehat{\bm{h}_{0}}$, and $\widehat{\bm{H}}_{i-1,\cdots,0}$ means $\widehat{\bm{H}_{i-1}},\widehat{\bm{H}_{i-2}},\cdots,\widehat{\bm{H}_{0}}$. The pair $(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})\in J_{\text{t}(d,\varepsilon)}^{(n)}$ can also be described as $\bm{m}_{i}^{n}$ and $\widehat{\bm{m}_{i}}^{n}$ are temporal distortion typical.
Lemma 2. If $(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})\in J_{\text{t}(d,\varepsilon)}^{(n)}$, then
$p(\widehat{\bm{m}_{i}}^{n}|\bm{m}_{i}^{n})\leq p(\widehat{\bm{m}_{i}}^{n}|\widehat{\bm{h}}_{i-1,\cdots,0})2^{n[I(\widehat{\bm{M}_{i}};\bm{M}_{i})-I(\widehat{\bm{M}_{i}};\widehat{\bm{H}}_{i-1,\cdots,0})+3\varepsilon]},$ (15)
where $\varepsilon>0$.
Proof:
$\begin{split}p(\widehat{\bm{m}_{i}}^{n}|\bm{m}_{i}^{n})=&\frac{p(\widehat{\bm{m}_{i}}^{n},\bm{m}_{i}^{n})}{p(\bm{m}_{i}^{n})}\\ =&p(\widehat{\bm{m}_{i}}^{n}|\widehat{\bm{h}}_{i-1,\cdots,0})\frac{p(\widehat{\bm{m}_{i}}^{n},\bm{m}_{i}^{n})}{p(\widehat{\bm{m}_{i}}^{n}|\widehat{\bm{h}}_{i-1,\cdots,0})p(\bm{m}_{i}^{n})},\\ \leq&\frac{p(\widehat{\bm{m}_{i}}^{n}|\widehat{\bm{h}}_{i-1,\cdots,0})2^{-n[H(\widehat{\bm{M}_{i}},\bm{M}_{i})-\varepsilon]}}{2^{-n[H(\widehat{\bm{M}_{i}}|\widehat{\bm{H}}_{i-1,\cdots,0})+\varepsilon]}2^{-n[H(\bm{M}_{i})+\varepsilon]}},\\ =&p(\widehat{\bm{m}_{i}}^{n}|\widehat{\bm{h}}_{i-1,\cdots,0})2^{n[I(\widehat{\bm{M}_{i}};\bm{M}_{i})-I(\widehat{\bm{M}_{i}};\widehat{\bm{H}}_{i-1,\cdots,0})+3\varepsilon]}.\end{split}$ (16)
Theorem 1. In time series estimation, the temporal correlation of the data can reduce the upper bound of the number of potential estimation results, and thus reduce the upper bound of the estimation error.
Proof: In the estimation process examined in this paper, as illustrated in Fig. 3, the estimation result $\widehat{\bm{M}_{i}}$ is obtained through the observation of CSI $\bm{H}_{i}$. Because the intricate relationship between $\bm{M}_{i}$ and $\bm{H}_{i}$, which defies explicit mathematical expression, it is challenging to analyze the impact of temporal correlation on estimation accuracy from the perspective of signal estimation theory. Note that the process can be regarded as the transmission of ground truth from the physical world to the estimator, which can be equivalent to a communication process. In this equivalent communication process, an asymptotically optimal communication mode is considered, namely random coding and joint typical decoding. The impact of temporal correlation on this communication mode is examined. The process of this communication mode is as follows:
  • Use block encoding, where each block consists of $n$ symbols, thereby encoding $\bm{m}_{i}^{n}$ in a single operation. Randomly generate a codebook $\mathcal{C}$ consisting $2^{nR}$ of sequences $\widehat{\bm{m}}^{n}$ drawn i.i.d. $\sim\prod_{j=1}^{n}p(\widehat{\bm{m}}^{(j)})$, where $R$ is the average amount of information in a single symbol. Index these codewords by $w\in\{1,2,\ldots,2^{nR}\}$. Reveal this codebook to the encoder and decoder.
  • Encode $\bm{m}_{i}^{n}$ by $w$ if there exists a $w$ such that $(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})$ is in the distortion typical set (or temporal distortion typical set for time series estimation). If there is more than one such $w$, send the least. If there is no such $w$, let $w=1$.
  • Decoding. The reproduced sequence is $c(w)=\widehat{\bm{m}_{i}}^{n}$.
Then, we can divide $\bm{m}^{n}\in\mathcal{M}^{n}$ into two categories:
  1. 1.
    $\bm{m}^{n}$ such that there exists a codeword $w$ that is distortion typical (or temporal distortion typical for time series estimation) with $\bm{m}^{n}$. In this category, let $E[d(\bm{M},\widehat{\bm{M}})]\leq D$, then $d(\bm{m}^{n},\widehat{\bm{m}}^{n})\leq D+\varepsilon$. The set of $\bm{m}^{n}$ that conforms to this category is referred to as the set of valid inputs for $c(\cdot)$, denoted by $V(c)=\{\bm{m}^{n}:\exists w\ \mathrm{with}\ (\bm{m}^{n},c(w))\in J_{(d,\varepsilon)}^{(n)}\}$ (or $V_{\text{t}}(c)=\{\bm{m}^{n}:\exists w\ \mathrm{with}\ (\bm{m}^{n},c(w))\in J_{\text{t}(d,\varepsilon)}^{(n)}\}$ for time series estimation).
  2. 2.
    $\bm{m}^{n}$ such that there does not exist a $w$ that is distortion typical (or temporal distortion typical for time series estimation) with $\bm{m}^{n}$. The distortion for any sequence is bounded by $d_{\text{max}}$. Let $P_{\text{e}}$ be the total probability of this category.
Then, the total distortion can be bounded by
$E[d(\bm{m}^{n},\widehat{\bm{m}}^{n})]\leq D+\varepsilon+P_{\text{e}}d_{\text{max}},$ (17)
Consider the first category. According to Lemma 1 and 2, we have
[…]
Since $I(\widehat{\bm{M}_{i}};\widehat{\bm{H}}_{i-1,\cdots,0})\geq 0$,
$\begin{split}p(\widehat{\bm{m}_{i}}^{n}|\bm{m}_{i}^{n})&2^{-n[I(\widehat{\bm{M}_{i}};\bm{M}_{i})+3\varepsilon]}\\ \leq&p(\widehat{\bm{m}_{i}}^{n}|\bm{m}_{i}^{n})2^{-n[I(\widehat{\bm{M}_{i}};\bm{M}_{i})-I(\widehat{\bm{M}_{i}};\widehat{\bm{H}}_{i-1,\cdots,0})+3\varepsilon]}.\end{split}$ (19)
From (19), the lower bound of $p(\widehat{\bm{m}_{i}}^{n})$ is less than that of $p(\widehat{\bm{m}_{i}}^{n}|\widehat{\bm{h}}_{i-1,\cdots,0})$. Since
$\sum_{\widehat{\bm{m}_{i}}^{n}\in{\widehat{\mathcal{M}}}^{n}}p(\widehat{\bm{m}_{i}}^{n})=\sum_{\widehat{\bm{m}_{i}}^{n}\in{\widehat{\mathcal{M}}}^{n}}p(\widehat{\bm{m}_{i}}^{n}|\widehat{\bm{h}}_{i-1,\cdots,0})=1,$ (20)
then the upper bound of the number of potential $\widehat{\bm{m}_{i}}^{n}$ exceeds that of $\widehat{\bm{m}_{i}}^{n}|\widehat{\bm{h}}_{i-1,\cdots,0}$, i.e., temporal correlation reduces the upper bound of the number of potential estimation results.
Consider the second category. In the absence of temporal correlation, $P_{\text{e}}$ can be represented as
$\begin{split}P_{\text{e}}&=\sum_{c}p(c)\sum_{\bm{m}_{i}^{n}:\bm{m}_{i}^{n}\notin V(c)}p(\bm{m}_{i}^{n}),\\ &=\sum_{\bm{m}_{i}^{n}}p(\bm{m}_{i}^{n})\sum_{c:\bm{m}_{i}^{n}\notin V(c)}p(c),\end{split}$ (21)
Define
[…]
Then,
$\begin{split}\sum_{c:\bm{m}_{i}^{n}\notin V(c)}p(c)&=\prod_{w}\{1-Pr[(\bm{m}_{i}^{n},c(w))\in J_{(d,\varepsilon)}^{(n)}]\},\\ &=[1-\sum_{\widehat{\bm{m}_{i}}^{n}}p(\widehat{\bm{m}_{i}}^{n})K(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})]^{nR}.\end{split}$ (23)
From Lemma 1,
$\begin{split}P_{\text{e}}\leq&\sum_{\bm{m}_{i}^{n}}p(\bm{m}_{i}^{n})\\ \times&\{1-\sum_{\widehat{\bm{m}_{i}}^{n}}p(\widehat{\bm{m}_{i}}^{n}|\bm{m}_{i}^{n})2^{-n[I(\widehat{\bm{M}_{i}};\bm{M}_{i})+3\varepsilon]}K(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})\}^{nR},\\ \leq&1-\sum_{\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n}}p(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})K(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})\\ &+\exp\{-2^{-n[I(\widehat{\bm{M}_{i}};\bm{M}_{i})+3\varepsilon]}2^{nR}\},\\ =&\text{Pr}[(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})\notin J_{(d,\varepsilon)}^{(n)}]+\exp\{-2^{n[R-I(\widehat{\bm{M}_{i}};\bm{M}_{i})-3\varepsilon]}\},\\ \leq&\varepsilon+\exp\{-2^{n[R-I(\widehat{\bm{M}_{i}};\bm{M}_{i})-3\varepsilon]}\}.\end{split}$ (24)
Thus,
$\begin{split}E&[d(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})]\\ &\leq D+d_{\text{max}}\exp\{-2^{n[R-I(\widehat{\bm{M}_{i}};\bm{M}_{i})-3\varepsilon]}\}+(1+d_{\text{max}})\varepsilon.\end{split}$ (25)
In the presence of temporal correlation, $P_{\text{e}}$ can be represented as
$\begin{split}P_{\text{e}}=&\sum_{c}p(c|\widehat{\bm{h}}_{i-1,\cdots,0})\sum_{\bm{m}_{i}^{n}:\bm{m}_{i}^{n}\notin V_{\text{t}}(c)}p(\bm{m}_{i}^{n}),\\ =&\sum_{\bm{m}_{i}^{n}}p(\bm{m}_{i}^{n})\sum_{c:\bm{m}_{i}^{n}\notin V_{\text{t}}(c)}p(c|\widehat{\bm{h}}_{i-1,\cdots,0}),\end{split}$ (26)
Then,
$\begin{split}\sum_{c:\bm{m}_{i}^{n}\notin V_{\text{t}}(c)}p&(c|\widehat{\bm{h}}_{i-1,\cdots,0})=\prod_{w}\{1-Pr[(\bm{m}_{i}^{n},c(w))\in J_{\text{t}(d,\varepsilon)}^{(n)}]\},\\ =&[1-\sum_{\widehat{\bm{m}_{i}}^{n}}p(\widehat{\bm{m}_{i}}^{n}|\widehat{\bm{h}}_{i-1,\cdots,0})K_{\text{t}}(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})]^{nR},\end{split}$ (27)
where
[…]
From Lemma 2,
$\begin{split}P_{\text{e}}\leq&1-\sum_{\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n}}p(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})K_{\text{t}}(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})\\ &+\exp\{-2^{-n[I(\widehat{\bm{M}_{i}};\bm{M}_{i})-I(\widehat{\bm{M}_{i}};\widehat{\bm{H}}_{i-1,\cdots,0})+3\varepsilon]}2^{nR}\},\\ \leq&\varepsilon+\exp\{-2^{n[R-I(\widehat{\bm{M}_{i}};\bm{M}_{i})+I(\widehat{\bm{M}_{i}};\widehat{\bm{H}}_{i-1,\cdots,0})-3\varepsilon]}\}.\end{split}$ (29)
Then, the total distortion should be bounded by
$\begin{split}E&[d(\bm{m}_{i}^{n},\widehat{\bm{m}_{i}}^{n})]\\ \leq&D+d_{\text{max}}\exp\{-2^{n[R-I(\widehat{\bm{M}_{i}};\bm{M}_{i})+I(\widehat{\bm{M}_{i}};\widehat{\bm{H}}_{i-1,\cdots,0})-3\varepsilon]}\}\\ &+(1+d_{\text{max}})\varepsilon.\end{split}$ (30)
Since
$\begin{split}D&+d_{\text{max}}\exp\{-2^{n[R-I(\widehat{\bm{M}_{i}};\bm{M}_{i})+I(\widehat{\bm{M}_{i}};\widehat{\bm{H}}_{i-1,\cdots,0})-3\varepsilon]}\}+(1+d_{\text{max}})\varepsilon\\ &\leq D+d_{\text{max}}\exp\{-2^{n[R-I(\widehat{\bm{M}_{i}};\bm{M}_{i})-3\varepsilon]}\}+(1+d_{\text{max}})\varepsilon,\end{split}$ (31)
the upper bound of the expected distortion is reduced. If the distortion function is defined as the estimation error, it can be argued that temporal correlation reduces the upper bound of the estimation error by constraining the upper bound of the number of plausible estimation outcomes. Notably, from (31), this reduction in the error bound stems from the mutual information term $I(\widehat{\bm{M}}_{i};\widehat{\bm{H}}_{i-1,\cdots,0})$, which quantifies the dependence between past CSI measurements and the current estimation result. If such dependence exists, then $I(\widehat{\bm{M}}_{i};\widehat{\bm{H}}_{i-1,\cdots,0})>0$, leading to a tighter upper bound on the error. Conversely, if no temporal correlation exists, $I(\widehat{\bm{M}}_{i};\widehat{\bm{H}}_{i-1,\cdots,0})=0$, and the error bound remains unchanged.
For HPE systems, a typical example is that it is often difficult for the model to distinguish the subject’s orientation (e.g. facing or back to the receiver), but if the subject has previously walked, the orientation can be determined based on the walking direction.
Figure 4: Schematic diagram of the AI-based real-time Wi-Fi sensing system.
III AI-based Real-time Wi-Fi Sensing and Visualization System
Building on our theoretical insights, we developed an AI-based real-time Wi-Fi sensing and visualization system. In this section, we provide a detailed description of the system’s design.
III-A System Overview
The proposed AI-based real-time Wi-Fi sensing and visualization system is illustrated in Fig. 4, which estimates human pose and location from Channel State Information (CSI) extracted from Wi-Fi signals. First, CSI data is collected in an indoor environment using the Linux 802.11n CSI Tool [18]. The raw data then undergoes preprocessing to remove static components and correct phase distortions. Following this, the processed data is fed into a neural network designed based on our theory, which outputs localization and human pose estimation (HPE) results. Finally, the results are plotted and displayed in real time. Note that the location of a person is defined as the center of the torso, while the pose is represented by the coordinates of each joint relative to this location.
III-B CSI Data
CSI characterizes how signals propagate from the transmitter to the receiver through multiple paths [16]. If a signal is transmitted through $L_{(n,m)}$ different propagation paths, the CSI characterizing the channel can be represented as
$h_{(n,m)}(f,t)=e^{-j\phi(f,t)}\sum_{l=1}^{L_{(n,m)}}\alpha_{l(n,m)}(f,t)e^{-j2\pi f\tau_{l(n,m)}(t)},$ (32)
where the subscript $(n,m)$ denotes the $n$-th receive antenna and the $m$-th transmit antenna, and $\tau_{l(n,m)}(t)$ is the propagation delay. If there are $N_{\text{t}}$ transmitting antennas and $N_{\text{r}}$ receiving antennas, the CSI can be represented as
$\bm{h}(f,t)=\begin{bmatrix}h_{(1,1)}(f,t)&\cdots&h_{(1,N_{\text{t}})}(f,t)\\ \vdots&\ddots&\vdots\\ h_{(N_{\text{r}},1)}(f,t)&\cdots&h_{(N_{\text{r}},N_{\text{t}})}(f,t)\\ \end{bmatrix}$. (33)
Notably, the influence of $\phi$ is highly stochastic, and in indoor environments, reflections from the target are often overwhelmed by those from walls. Both factors can significantly degrade the model performance. Therefore, preprocessing of the CSI data is essential to mitigate these adverse effects.
III-C CSI Preprocessing
The preprocessing method employed in this work follows the framework introduced in [9], leveraging the CSI from a single receiving antenna as a reference to reconstruct the dynamic components present in the CSI from all other antennas. However, since the aforementioned study utilized systems with a single transmitting antenna and multiple receiving antennas, our subsequent derivation in this section demonstrates that directly applying this method to a more general system with $N_{\text{t}}$ transmitting antennas would result in discarding $N_{\text{t}}-1$ elements of the CSI matrix. To address this limitation, we have introduced necessary modifications to the method.
We begin by reshaping the original CSI matrix $\bm{h}(f,t)$ into an $N_{\text{t}}N_{\text{r}}\times 1$ column vector $\bm{h}^{\prime}(f,t)$. We have
$\bm{h}^{\prime}(f,t)=$ $\left[h_{(1,1)}(f,t)\cdots h_{(1,N_{t})}(f,t),\right$. (34)
$\left.h_{(2,1)}(f,t)\cdots h_{(2,N_{t})}(f,t)\cdots h_{(N_{\text{r}},N_{\text{t}})}(f,t)\right]^{T},$
Treat $\bm{h}^{\prime}(f,t)$ as being composed of both $\bm{h}_{\text{s}}(f)$, which is contributed solely by paths affected by static objects, and $\bm{h}_{\text{d}}(f,t)$, which is contributed by paths influenced by moving objects, i.e.,
$\bm{h}^{\prime}(f,t)=(\bm{h}_{\text{s}}(f)+\bm{h}_{\text{d}}(f,t))e^{-j\phi(f,t)}.$ (35)
Then we have
$\begin{split}\bm{h}^{\prime}(f,t)\bm{h}^{\prime\dagger}(f,t)=&\bm{h}_{\text{s}}(f)\bm{h}_{\text{s}}^{\dagger}(f)+\bm{h}_{\text{s}}(f)\bm{h}_{\text{d}}^{\dagger}(f,t)\\ &+\bm{h}_{\text{d}}(f,t)\bm{h}_{\text{s}}^{\dagger}(f)+\bm{h}_{\text{d}}(f,t)\bm{h}_{\text{d}}^{\dagger}(f,t),\end{split}$ (36)
where $\dagger$ means conjugate transpose. Let the element of matrix $\bm{A}$ in row n and column m be $\bm{A}_{(n,m)}$, and the $n\text{th}$ element of vector $\bm{B}$ be $\bm{B}_{(n)}$. In indoor environments, the reflected signals of walls are very strong, so $\forall n\in\left(0,N_{\text{t}}N_{\text{r}}\right]$, $\exists\left\lVert\bm{h}_{\text{s}(n)}\right\rVert\gg\left\lVert\bm{h}_{\text{d}(n)}\right\rVert$. Then we have
[…]
where $0<n\leq N_{\text{r}},0<m\leq N_{\text{t}}$. So
$\bm{h}^{\prime}(f,t)\bm{h}^{\prime\dagger}(f,t)\approx\bm{h}_{\text{s}}(f)\bm{h}_{\text{s}}^{\dagger}(f)+\bm{h}_{\text{s}}(f)\bm{h}_{\text{d}}^{\dagger}(f,t)+\bm{h}_{\text{d}}(f,t)\bm{h}_{\text{s}}^{\dagger}(f),$ (38)
where $\bm{h}_{\text{s}}(f)\bm{h}_{\text{s}}^{\dagger}(f)$ can be eliminated by a high-pass filter (HPF), i.e.,
$\begin{split}\bm{D}(f,t)&=\text{HPF}(\bm{h}^{\prime}(f,t)\bm{h}^{\prime\dagger}(f,t))\\ &\approx\bm{h}_{\text{s}}(f)\bm{h}_{\text{d}}^{\dagger}(f,t)+\bm{h}_{\text{d}}(f,t)\bm{h}_{\text{s}}^{\dagger}(f).\end{split}$ (39)
For $\bm{h}_{\text{d}}(f,t)$, which we care about, we can’t get it independently from (39).
Take $N_{\text{ref}}$ elements with the largest average module in $\bm{h}^{\prime}(f,t)$ for reference, $1\leq N_{\text{ref}}<N_{\text{t}}N_{\text{r}}$. For convenience, the reference elements are assumed to be the initial $N_{\text{ref}}$ elements in $\bm{h}^{\prime}(f,t)$. Since the CSI estimation algorithm estimates from large to small, the position of reference elements are actually consistent with the assumption. Then we have
$\bm{h}^{\prime}(f,t)+\binom{\bm{\beta}}{\bm{0}}=\left[(\bm{h}_{\text{s}}(f)+\binom{\bm{\beta}}{\bm{0}}e^{j\phi(f,t)})+\bm{h}_{\text{d}}(f,t)\right]e^{-j\phi(f,t)},$ (40)
where $\bm{\beta}$ is a vector with size $N_{\text{ref}}\times 1$. Then
$\begin{split}&\left(\bm{h}^{\prime}(f,t)+\binom{\bm{\beta}}{\bm{0}}\right)\left(\bm{h}^{\prime}(f,t)+\binom{\bm{\beta}}{\bm{0}}\right)^{\dagger}\\ &=\left(\bm{h}_{\text{s}}(f)+\binom{\bm{\beta}}{\bm{0}}e^{j\phi(f,t)}\right)\left(\bm{h}_{\text{s}}(f)+\binom{\bm{\beta}}{\bm{0}}e^{j\phi(f,t)}\right)^{\dagger}\\ &+\left(\bm{h}_{\text{s}}(f)+\binom{\bm{\beta}}{\bm{0}}e^{j\phi(f,t)}\right)\bm{h}_{\text{d}}^{\dagger}(f,t)\\ &+\bm{h}_{\text{d}}(f,t)\left(\bm{h}_{\text{s}}(f)+\binom{\bm{\beta}}{\bm{0}}e^{j\phi(f,t)}\right)^{\dagger}\\ &+\bm{h}_{\text{d}}(f,t)\bm{h}_{\text{d}}^{\dagger}(f,t).\end{split}$ (41)
Then,
$\begin{split}\bm{D}^{\prime}(f,t)=&\text{HPF}\left[\left(\bm{h}^{\prime}(f,t)+\binom{\bm{\beta}}{\bm{0}}\right)\left(\bm{h}^{\prime}(f,t)+\binom{\bm{\beta}}{\bm{0}}\right)^{\dagger}\right]\\ =&\left(\bm{h}_{\text{s}}(f)+\binom{\bm{\beta}}{\bm{0}}e^{j\phi(f,t)}\right)\bm{h}_{\text{d}}^{\dagger}(f,t)\\ &+\bm{h}_{\text{d}}(f,t)\left(\bm{h}_{\text{s}}(f)+\binom{\bm{\beta}}{\bm{0}}e^{j\phi(f,t)}\right)^{\dagger}\\ &+\bm{h}_{\text{d}}(f,t)\bm{h}_{\text{d}}^{\dagger}(f,t).\end{split}$ (42)
Assume $\left\lvert\bm{\beta}_{(m)}\right\rvert\gg\left\lvert\bm{h}_{\text{s}(n)}(f)\right\rvert$, then we have
$\bm{D}^{\prime}_{(n,m)}(f,t)\approx(\bm{h}_{\text{s}(m)}(f)+\bm{\beta}_{(m)}e^{j\phi(f,t)})^{\ast}\bm{h}_{\text{d}(n)}(f,t),$ (43)
where $1\leq m\leq N_{\text{ref}}<n\leq N_{\text{t}}N_{\text{r}}$, and $\ast$ means conjugate. Let $\bm{\beta}_{(m)}$ the same phase as $\bm{h}_{(m)}(f,t)$, $1\leq m\leq N_{\text{ref}}$. Since $\left\lVert\bm{h}_{\text{s}(l)}\right\rVert\gg\left\lVert\bm{h}_{\text{d}(l)}\right\rVert$, $0<l\leq N_{\text{t}}N_{\text{r}}$, then $\frac{\bm{\beta}_{(m)}e^{j\phi(f,t)}}{\left\lVert\bm{\beta}_{(m)}e^{j\phi(f,t)}\right\rVert}\approx\frac{\bm{h}_{\text{s}(m)}(f)}{\left\lVert\bm{h}_{\text{s}(m)}(f)\right\rVert}$. At this time, the effect of $\bm{\beta}_{(m)}$ on $\bm{h}_{\text{s}(m)}(f)$ can be considered as increasing the modulus of $\bm{h}_{\text{s}(m)}(f)$, and the phase change of $\bm{h}_{\text{s}(m)}(f)$ is small. Let $\bm{h}^{\prime}_{\text{s}(m)}(f)=\bm{h}_{\text{s}(m)}(f)+\bm{\beta}_{(m)}e^{j\phi(f,t)}$, then if $1\leq m\leq N_{\text{ref}}<n\leq N_{\text{t}}N_{\text{r}}$,
$\bm{D}^{\prime}_{(n,m)}(f,t)\approx\bm{h}^{\prime\,\ast}_{\text{s}(m)}(f)\bm{h}_{\text{d}(n)}(f,t),$ (44)
According to (44), when $1\leq n\leq N_{\text{ref}}<m\leq N_{\text{t}}N_{\text{r}}$, $\bm{D}^{\prime}_{(n,m)}(f,t)$ is linear to the dynamic component. At this time, in the original $N_{\text{t}}N_{\text{r}}\times 1$ CSI vector $\bm{h}^{\prime}(f,t)$, we obtain $N_{\text{t}}N_{\text{r}}-N_{\text{ref}}$ elements of $\bm{D}^{\prime}_{(n,m)}(f,t)$ that are approximately linear to the dynamic component at the cost of abandoning $N_{\text{ref}}$ reference elements. Since the dynamic component is critical for HPE, we aim to maximize the number of dynamic components that are linear with respect to $\bm{D}^{\prime}_{(n,m)}(f,t)$. To achieve this, we set $N_{\text{ref}}=1$.
III-D Network Design
The architecture of the proposed network (Fig. 5) consists of an encoder for feature extraction and data compression, followed by a neck for intermediate processing, and a head for final pose estimation.
We employ an encoder to compress the input data and perform feature extraction. We began by training an autoencoder, and the output of its encoder portion was then used as the input to the subsequent components of the network. In this encoder, we combine Convolutional Neural Networks (CNN) and MaxPooling layers to perform feature extraction and compression on the input CSI data. To enhance the encoder’s sensitivity to temporal dependencies in the input data, we also incorporated two layers of Convolutional Gate Recurrent Units (ConvGRU). Furthermore, to investigate the impact of temporal correlation, the implemented ConvGRU architecture includes a hyperparameter that controls the maximum length of processable sequential data.
Figure 5: Illustration of the network structure, where we incorporate a hyperparameter in the stacked ConvGRU layers to control the maximum processable sequence length. The depth head and 2D pose estimation head work jointly to achieve 3D human pose estimation.
We employ an Hourglass network as the neck to further extract complex features embedded in the encoded CSI data. The network is equipped with three dedicated heads: one for 2D pose estimation, another for depth estimation, and the third for localization. The 2D pose estimation head and the depth estimation head jointly form the 3D HPE capability by integrating their respective outputs. To accelerate model convergence, we follow a common practice in computer vision by having each head output heatmaps instead of directly predicting the coordinates.
III-E Loss Function
The loss function is composed of a weighted sum of the individual loss functions from each head, i.e.,
$\mathcal{L}=a\left[b\mathcal{L}_{\text{Depth}}+\left(1-b\right)\mathcal{L}_{\text{HPE}}\right]+\left(1-a\right)\mathcal{L}_{\text{Location}},$ (45)
where $\mathcal{L}_{\text{head}}$ denote the loss function of a head, with weights $a$ and $b$ satisfying $a,b\in\left(0,1\right)$.
The loss function for each head comprises two components: a peak location error and a pixel-wise error. To combine the errors of these two quantities with different physical dimensions, we used the normalized mean square error (NMSE), which is a dimensionless measure. We denote the output of a head as $\widehat{\bm{m}}_{\text{head}}$ and its corresponding ground truth as $\bm{m}_{\text{head}}$. Thus, the error for each head can be expressed as
$\mathcal{L}_{\text{Depth}}=\frac{\text{MSE}\left(\text{argmax}\left(\widehat{\bm{m}}_{\text{Depth}}\right)\right)}{\sigma^{2}\left(\text{argmax}\left(\bm{m}_{\text{Depth}}\right)\right)}c+\frac{\text{MSE}\left(\widehat{\bm{m}}_{\text{Depth}}\right)}{\sigma^{2}\left(\bm{m}_{\text{Depth}}\right)}\left(1-c\right),$ (46)
$\mathcal{L}_{\text{HPE}}=\frac{\text{MSE}\left(\text{argmax}\left(\widehat{\bm{m}}_{\text{HPE}}\right)\right)}{\sigma^{2}\left(\text{argmax}\left(\bm{m}_{\text{HPE}}\right)\right)}c+\frac{\text{MSE}\left(\widehat{\bm{m}}_{\text{HPE}}\right)}{\sigma^{2}\left(\bm{m}_{\text{HPE}}\right)}\left(1-c\right),$ (47)
$\begin{split}\mathcal{L}_{\text{Location}}=&\frac{\text{MSE}\left(\text{argmax}\left(\widehat{\bm{m}}_{\text{Location}}\right)\right)}{\sigma^{2}\left(\text{argmax}\left(\bm{m}_{\text{Location}}\right)\right)}c\\ &+\frac{\text{MSE}\left(\widehat{\bm{m}}_{\text{Location}}\right)}{\sigma^{2}\left(\bm{m}_{\text{Location}}\right)}\left(1-c\right),\end{split}$ (48)
where $\mathrm{MSE}(\cdot)$ denotes the mean squared error, $\sigma^{2}(\cdot)$ represents the variance, $\mathrm{argmax}(\cdot)$ indicates the operation of finding the position of the maximum value for each channel, and $c$ is a weighting coefficient satisfying $c\in[0,1]$, which can be adjusted based on the specific objectives at different stages of training. During the initial training stages, a lower value is assigned to $c$ to facilitate rapid convergence. In later training, $c$ is increased to refine the accuracy of peak localization.
III-F Real-Time Data Processing Design
To meet real-time requirements, the following three issues must be addressed:
  • Receiver system limitation. The Linux 802.11n CSI Tool in the Wi-Fi sensing system operates on Ubuntu 14.04, which lacks support for running neural networks.
  • Inefficient reading of numerous small files. The CSI acquisition is implemented using the Linux 802.11n CSI Tool, which is written in C, whereas neural networks typically rely on Python. Thus, CSI data cannot be passed directly via memory; it must first be written to disk and then read by Python. Although CSI data is collected and saved at a rate of 1000 fps, reading it from the hard disk (HDD or SSD) is far slower, thus severely limiting the subsequent data processing frequency.
  • High overhead of scanning filenames. Reading CSI with Python requires scanning the entire directory to obtain all filenames before loading the oldest unread file to maintain sequential order. This repeated directory scanning further slows down the process.
Since the neural network structure is lightweight enough, its computational overhead is not the bottleneck for real-time performance. To resolve the above issues, we built a real-time sensing and visualization system illustrated in Fig. 6.
Figure 6: Data processing flowchart of the real-time sensing and visualization system. The receiver streams the captured CSI data in real time to a host computer. All subsequent steps — including data reading, processing, and result visualization — are integrated into a processor on the host computer.
To address the limitations of the receiver system, the receiver streams the captured CSI data in real time to a host computer. The host computer is a commodity portable personal computer, an HP OMEN 8 Pro, equipped with an NVIDIA GeForce RTX 3060 graphics card. All subsequent steps — including data reading, processing, and result visualization — are integrated into a processor on the host computer. To eliminate the overhead of scanning file names, we set up an inotify-based watcher. Whenever a new file is created, inotify generates a file update event, from which the file name can be directly obtained without scanning the entire directory. To address the inefficiency of reading a large number of small files, we adopted a multi-process and multi-thread parallel reading and processing scheme and employed queues to manage data flow between different processes. Data is pushed to the tail of the queue upon arrival, and the system pulls (and removes) data from the head of the queue when needed. This push/pull mechanism ensures that the order of data processing remains consistent with the original sequence in which the receiver collects the CSI data.
IV Experimental Setup and Results
IV-A CSI Collection
The CSI data was collected in $45$ distinct indoor environments. To improve data collection efficiency, we simultaneously used two receivers placed at different locations during each data acquisition session, along with one transmitter, resulting in a total of three personal computers employed in the system, each equipped with an Intel 5300 NIC and configured with the Linux 802.11n CSI Tool [18]. The two receivers were synchronized using the network time protocol (NTP), achieving an average synchronization error of less than $10\text{ ms}$. The distance between the transmitter and one of the receivers was $7\text{ m}$. Both the transmitter and the receivers were equipped with three antennas each, and the receiving antennas were directly facing the transmitting antennas. The system measured CSI over $30$ subcarriers in the $5320\text{ MHz}$ band with a bandwidth of $10\text{ MHz}$, at a rate of $1000$ packets per second. The dimension of the CSI obtained per packet is $3\times 3\times 30$. Following the approach in [9], we preprocess the CSI by retaining the reference signal, resulting in a signal of dimension $9\times 30$. Since this signal is complex-valued, both its amplitude and phase are separately fed into the network. Thus, the final dimension of the preprocessed CSI is $2\times 9\times 30$ per packet.
Figure 7: Average Estimation Error over Time. (a) Average pose estimation error at different timestamps. (b) Average localization error at different timestamps.
IV-B 3D Human Pose Collection
We developed a binocular stereo vision-based system to capture 3D human pose data as ground truth. This system ensures that the coordinates of the major joints in the reconstructed human skeleton maintain high spatial consistency with their corresponding real-world coordinates.
A total of six participants with varying body shapes, heights, and genders were involved in the data collection. Each participant performed four types of actions: walking, horizontally raising the left arm, horizontally raising the right arm, and horizontally raising both arms. The sample size for each action is balanced. During walking, the participants swung their arms naturally back and forth at their sides. When raising arms horizontally, the arms were kept within the same plane as the torso.
The system employs an Orbbec Gemini 336L structured light stereo camera with a field of view (FoV) of $90^{\circ}\text{(horizontal)}\times 65^{\circ}\text{(vertical)}$, a resolution of $1280\times 800$ pixels, and a frame rate of $30\text{ fps}$. The camera was placed adjacent to one receiver, facing the same direction as the receiver’s antennas.
Collected video frames are processed in two stages:
  • In the first stage, a 3D multi-person pose estimation (3DMPPE) method [20] is used to extract the 3D human skeletal structure. We selected 16 major joints that represent the essential human pose information.
  • In the second stage, the Semi-Global Block Matching (SGBM) algorithm [21] is applied to compute the depth map of the current frame. Then, the YOLOv10 model [22] detects the pixel coordinates of the human torso center in the frame. These coordinates are used to index the depth map, retrieving the world coordinates (i.e., coordinates relative to the camera) of the torso center.
Using this torso center location, the 3D human pose obtained in the first stage is adjusted to derive the joint coordinates in real physical space. The system achieves centimeter-level average error between the captured joint positions and their ground-truth locations.
Since two receivers were employed to capture CSI while only one camera was available, the camera was co-located with one of the receivers. Consequently, the captured joint coordinates are expressed in that receiver’s coordinate frame. To recover the human pose from the perspective of the other receiver, the acquired pose data must undergo a rigid transformation consisting of translation and rotation.
IV-C Data Alignment and Segmentation
The captured human pose data and the video frames share the same frame rate of $30\text{ Hz}$, whereas the CSI is sampled at $1000\text{ Hz}$. Temporal alignment between the pose data and CSI is achieved using timestamps. For each pose instance, we assign the $32$ temporally nearest CSI frames, thereby forming a data sample comprising a CSI tensor of dimension $32\times 2\times 9\times 30$ as input to the network, with the corresponding human pose serving as the ground truth.
A total of $24208$ data samples were collected. The dataset was partitioned into $18768$ samples for training, $3264$ for validation, and $2176$ for testing. To evaluate the influence of temporal correlation, the data were arranged chronologically and divided into contiguous blocks. In the collected dataset, the length of continuous data sequences ranges from a minimum of 68 to a maximum of 90. We set the block length to 17 (i.e., one-fourth of 68). After segmentation, data shorter than one block length are discarded. During both training and testing, a batch size of $16$ such blocks was utilized.
IV-D Experimental Results
To validate Theorem 1, we trained models under two settings: one without temporal memory and another with temporal memory, resulting in two sets of model parameters denoted as HPE_I and HPE_T, respectively. As shown in Fig. 7, the errors of HPE_T are consistently lower than those of HPE_I for both localization and pose estimation tasks. Furthermore, the errors of HPE_T exhibit a decreasing trend over time, confirming the performance gains brought by temporal correlation.
Figure 8: Example of temporal correlation gain, where the blue skeletons represent the ground truth, while the red skeletons indicate the estimation results. It can be observed that the pose estimation results become progressively more accurate over time.
We use the example in Fig. 8 to illustrate how temporal correlation reduces estimation errors by narrowing down the number of plausible estimation outcomes. In the figure, the blue skeletons represent the ground truth, while the red skeletons indicate the estimation results. At timestamp 1, the model infers that the subject is raising his hand, and the estimated orientation deviates significantly from the ground truth. By timestamp 2, the estimated pose shows higher similarity to the ground truth, but the orientation remains inaccurate. By timestamp 6, both the pose and orientation converge closely to the ground truth.
Figure 9: (a) Loss curve under direct training, where HPE_T performs slightly inferior to HPE_I. (b) Loss curve after pre-training, where HPE_T significantly outperforms HPE_I, indirectly indicating the performance gain brought by prior information.
Moreover, during training we observed that when trained directly, the loss of HPE_T exhibited pronounced oscillations, and its performance was even slightly inferior to that of HPE_I, as shown in Fig. 9a. However, when HPE_I was first trained and its weights were used as pre-trained initialization for HPE_T, the loss of HPE_T became more stable compared to direct training, and its performance showed a significant improvement over HPE_I, as demonstrated in Fig. 9b. This suggests that allowing the model to first fully learn the structural information of the human body before incorporating temporal correlation leads to superior performance, providing indirect support for the performance gain brought by prior information. Furthermore, it shows that for tasks requiring performance improvement through temporal correlation, this pre-training approach can maximize the utilization of temporal correlation.
TABLE I: The mean absolute error (MAE) of HPE.
Joint Thorax Right Shoulder Right Elbow Right Wrist Left Shoulder Left Elbow Left Wrist
MAE (m) 0.0886 0.1565 0.3678 0.4763 0.1799 0.3517 0.4856
Joint Right Hip Right Knee Right Ankle Left Hip Left Knee Left Ankle Pelvis Head Average
MAE (m) 0.1513 0.1599 0.2197 0.1335 0.1393 0.1787 0.0872 0.1069 0.2189
Figure 10: (a) CDF of HPE_T for the human pose estimation task. (b) CDF of HPE_T for the indoor localization task.
[…]
V Conclusions
This paper theoretically analyzes and identifies two primary sources of performance gains that AI brings to Wi-Fi sensing under hardware-constrained conditions. Building on our theoretical findings, we develop an AI-based real-time Wi-Fi sensing and visualization system. We argue that the gains of AI primarily stem from the effective utilization of prior information and temporal correlation. The prior information enables AI to generate relatively detailed sensing results based on a vague perception of the target’s state by leveraging the structural priors learned during training, without the need to perceive every detail of the target meticulously. This allows AI-based sensing to surpass the system’s radar aperture. Meanwhile, temporal correlation enables AI to reduce the space of plausible sensing outcomes by leveraging the correlation across sequential data, thereby improving sensing accuracy.
The AI-based real-time Wi-Fi sensing and visualization system performs indoor localization and human pose estimation using only a single transceiver pair. The system achieved an average localization error of $0.6124\text{ m}$, an average pose estimation error of $0.2189\text{ m}$, and an average frame rate of at least $42\text{ fps}$ on commodity hardware. Experimental results confirm the performance gains brought by temporal correlation and provide indirect evidence for the benefits of prior information. Furthermore, we found that to fully leverage temporal correlation, the AI model must first undergo sufficient pre-training under a non-temporal setting to adequately learn prior structural knowledge, before being fine-tuned with temporal awareness. Otherwise, direct training tends to lead the model to converge to suboptimal local minima.